Amy Dyken, controller at Bridgeport Pharmaceutical Industries, a public company, is currently preparing the calculation for basic and diluted earnings per share and the related disclosure for Bridgeport’s financial statements. Below is selected financial information for the fiscal year ended June 30, 2020.

Bridgeport Pharmaceutical Industries
Selected Balance Sheet Information
June 30, 2020

Long-term debt
   Notes payable, 10%

$990,000

   9% convertible bonds payable

5,040,000

   10% bonds payable

6,080,000

     Total long-term debt

$12,110,000

Shareholders’ equity
   Preferred stock, 5% cumulative, $50 par value, 104,000 shares authorized, 26,000 shares issued and outstanding

$1,300,000

   Common stock, $1 par, 10,000,000 shares authorized, 1,000,000 shares issued and outstanding

1,000,000

   Additional paid-in capital

3,990,000

   Retained earnings

6,090,000

     Total shareholders’ equity

$12,380,000


The following transactions have also occurred at Bridgeport.

1. Options were granted on July 1, 2019, to purchase 180,000 shares at $15 per share. Although no options were exercised during fiscal year 2020, the average price per common share during fiscal year 2020 was $20 per share.
2. Each bond was issued at face value. The 9% convertible bonds will convert into common stock at 50 shares per $1,000 bond. The bonds are exercisable after 5 years and were issued in fiscal year 2019.
3. The preferred stock was issued in 2019.
4. There are no preferred dividends in arrears; however, preferred dividends were not declared in fiscal year 2020.
5. The 1,000,000 shares of common stock were outstanding for the entire 2020 fiscal year.
6. Net income for fiscal year 2020 was $1,490,000, and the average income tax rate is 20%.


For the fiscal year ended June 30, 2020, calculate the following for Bridgeport Pharmaceutical Industries. (Round answers to 2 decimal places, e.g. $2.45.)

(a) Basic earnings per share.

Basic earnings per share

$


(b) Diluted earnings per share.

Diluted earnings per share

$
Answer #1

Part A

Basic earnings per share = (Net Income - Preferred Dividends)/ Weighted-Average Shares = (1490000-65000)/ 1000000 = $1.43

Preferred dividend = 26000 preferred shares * $50/share * 5% = $65000

Therefore,

Basic earnings per share: $1.43

Part B

Test for dilution:

Convertible Bonds:

Interest on convertible bonds (net of tax)/ Converted common shares = (5040000*9%*(1-20%))/((5040000/1000)*50) =221452/271782= 1.44

As 1.44 is greater than the 1.43 Basic EPS, the convertible bonds are antidilutive

Stock Options

As Average market price ($20) > option price ($15), the stock options are dilutive.

Incremental Shares = Market Price - Option Price/Market Price * Number of options or warrants = ((20-15)/20)*180000 = 45000 shares

Diluted earnings per share = Net Income - Preferred Dividends + Interest (net of tax)/ Average Common Shares + Potentially Dilutive Common Shares = (1490000-65000)/(1000000+45000) = $1.36

Therefore,

Diluted earnings per share: $1.36